The capacitance of parallel plate capacitor is increased by a factor of 5 when a dielectric material fills the space between its plates.what is the relative permittivity of the dielectric material? If this material is placed in between the plates of a cylindrical capacitor of outer and inner raddi 12cm and 10 cm respectively, calculate the capacitance per unit length of the cylindrical capacitor.

Grade:12th pass

1 Answers

Eshan
5 years ago
The capacitance of a capacitor\proptorelative permittivity of the dielectric material.
C=\dfrac{\epsilon_r\epsilon_0 A}{d}
Hence the relative permittivity of the dielectric material is 5.

The capacitance of a cyclindrical capacitor is given by
C_{cylinder}=\dfrac{2\pi \epsilon_r\epsilon_0}{ln(\dfrac{r_2}{r_1})}
Hence the capacitance per unit length=\dfrac{2\pi (5)\epsilon_0}{ln(\dfrac{12}{10})}=1.5nF